OpenStudy (anonymous):

1. Jameston High School has 2020 students. Its enrollment has been decreasing at a rate of 125 students per year. Townsville High School has 1075 students. Its enrollment has been increasing at a rate of 80 students per year. Assuming the enrollment rates continue: A. Write an equation to model this. B. When will they have the same enrollment? C. What is that enrollment amount?

OpenStudy (kash_thesmartguy):

J=2020-125x T=1075+80x

OpenStudy (anonymous):

so im a bit confused why that ? sorry when it comes to math im completly lost

OpenStudy (kash_thesmartguy):

x represents years, -125 represents loss of students and J represents Jamestown School

OpenStudy (anonymous):

how can i write that into an answer for the question

OpenStudy (kash_thesmartguy):

@dan815

OpenStudy (anonymous):

the equations @Kash_TheSmartGuy put up are the answers for part A

OpenStudy (anonymous):

for B you need to set the equations equal to each other and solve for x 2020-125x=1075+80x

OpenStudy (kash_thesmartguy):

2020-125x=1075+80x Solve for x and there is the time it will take them to become the same.

OpenStudy (anonymous):

x=18941 ?

OpenStudy (anonymous):

189/41

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

for C plug 189/41 in for x in either of the equations from A to get the enrollment

OpenStudy (anonymous):

59195/41 this is what i got ?

OpenStudy (anonymous):

yes. you should probably write that as a decimal and round to the nearest whole. Can't have part of a student

OpenStudy (anonymous):

1443.7804878 im not so sure how to round tho

OpenStudy (kash_thesmartguy):

Do you have to round to the nearest hundreth, tenth or what?

OpenStudy (anonymous):

1444

OpenStudy (kash_thesmartguy):

1443.8 or 1443.78 or 1444

OpenStudy (kash_thesmartguy):

You get it, right? @Andrews_girl

OpenStudy (anonymous):

yes i understand

OpenStudy (kash_thesmartguy):

k.. good, any other questions?

OpenStudy (kash_thesmartguy):

Please end this question and then start a new question if you have one.

OpenStudy (anonymous):

yes actually oh and thank you for being patient with me , i must seem like such an airhead

OpenStudy (kash_thesmartguy):

Nope, it's good that you're eager to learn!

OpenStudy (kash_thesmartguy):

Learning is good and don't feel bad if you don't know something, someone's always around to help you out. Actually, this was a bit of a learning experience for me?

OpenStudy (anonymous):

well yeah i guess your right , i just hate being really bad at math.

OpenStudy (anonymous):

given : f(x)=8-4x A. Find f(=2) b. Find x when, f(x)=8 c. graph the function D. state the domian and range of the function e. use the graph to find x when f(x)=0. show the details on the graph f. Use the graph to find f(x) when x = 3. Show details on the graph